OpenStudy (anonymous):

How do i verify this 2tan(X)/(1+tan(X)^(2))=sin(2x)?

OpenStudy (anonymous):

\[\frac{2 \tan x}{1+\tan^2 x}=\frac{2\frac{\sin x}{\cos x}}{1+\frac{\sin^2 x}{\cos^2 x}}\] multiply num and denum by \(\cos^2 x\) and see what happens?